OpenStudy (anonymous):

what is the degrees for 11pi/10 radians and -2pi/3 radians

OpenStudy (anonymous):

(11*180)/10 = something degrees 360 -[(-2*180) /3] = something degrees

OpenStudy (anonymous):

\[\frac{\left(\frac{11 \pi }{10}\right)}{\pi }180=190{}^{\circ} \]\[\frac{\left(-\frac{2 \pi }{3}\right)}{\pi }180=-120 \]